# Simplified vine copula models: state of science and affairs

### v1

@@ Abstract
Vine copula models have become highly popular practical tools for modeling multivariate dependencies. To maintain tractability, a commonly employed simplifying assumption is that conditional copulas remain unchanged by the conditioning variables. This assumption has sparked a somewhat polarizing debate within our community. The fact that much of this dispute occurs outside the public record has placed the field in an unfortunate position, impeding scientific progress. In this article, I will review what we know about the flexibility and limitations of simplified vine copula models, explore the broader implications, and offer my own, hopefully reconciling, perspective on the issue.
@@ 2.2 The simplifying assumption in the bivariate case
Now consider the following conditional version of Sklar’s theorem (Patton,, 2001). For any index set $D\subset\{1,\ldots,d\}$, indices $i,j\notin D$, and every value $\bm{x}_{D}\in\mathds{R}^{|D|}$, the conditional distribution of $(X_{i},X_{j})$ given $\bm{X}_{D}=\bm{x}_{D}$ can be expressed as
@@
where $C_{i,j|D}(\cdot\mid\bm{x}_{D})$ is a bivariate copula for every value of $\bm{x}_{D}$. [^1] In the vine copula literature, the conditional copula is more commonly denoted by $C_{i,j;D}$ to distinguish it from the conditional distribution of $(U_{i},U_{j})$ given $\bm{U}_{D}$, which is not a copula. The latter distribution or related conditionals do not appear in this article, so there should be no confusion. More precisely, it is the conditional joint distribution functions of the random vector $(U_{i|D},U_{j|D})$ given $\bm{X}_{D}=\bm{x}_{D}$, where
@@
The Gaussian distribution provides an accessible analogy. Let $\bm{X}\sim\mathcal{N}(0,\Sigma)$ and define $\varepsilon_{i|D}$ and $\varepsilon_{j|D}$ as the residuals from regressing $X_{i}$ and $X_{j}$ on $\bm{X}_{D}$, scaled to unit variance. This is, in essence, what the operation $F_{i|D}(X_{i}\mid\bm{X}_{D})$ does. The conditional copula $C_{i,j|D}(\cdot\mid\bm{x}_{D})$ characterizes the dependence of the regression residuals $\varepsilon_{i|D}$ and $\varepsilon_{j|D}$. In the Gaussian world, dependence is synonymous with correlation, so we can think of $C_{i,j|D}(\cdot\mid\bm{x}_{D})$ as the conditional correlation coefficient $\rho_{i,j|D}(\bm{x}_{D})={\mathrm{Corr}}(\varepsilon_{i|D},\varepsilon_{j|D}\mid\bm{X}_{D}=\bm{x}_{D})$. The simplifying assumption asserts that this correlation coefficient does not change with the value $\bm{x}_{D}$. Indeed, the Gaussian distribution satisfies $\rho_{i,j|D}\equiv\rho_{i,j;D}$, where $\rho_{i,j;D}$ is the partial correlation coefficient of $(X_{i},X_{j})$ given $\bm{X}_{D}$. This equality does not typically hold outside of Gaussian models, however. For more details, we refer to Gijbels et al., (2015) and Spanhel and Kurz, (2016) who expand on this idea in terms of conditional and partial copulas.
@@ 2.3 The simplifying assumption in vine copulas
3 What we know about the limitations simplified models
@@ 3.5 Simplified vines as statistical models
Simplified vine copulas are statistical models for the dependence structure of a random vector. From a philosophical point of view, they are no different from other statistical models serving other purposes. Models are most often approximations of reality, and as such, mostly wrong. To paraphrase a famous quote by George Box: The question is not whether they are wrong, but whether they are useful.
@@
The second point emphasizes that we cannot disentangle the approximation qualities of a model from our ability to estimate it. Generally speaking, more flexible models are harder to estimate—the ubiquitous bias-variance trade-off. In extreme cases, the model can be so large that consistent estimation is impossible. This is the case, for example, if $\Theta$ is the collection of all probability densities on $\mathds{R}^{d}$. Parametric families of densities on the other hand can be estimated at $O_{p}(n^{-1/2})$ rate under mild conditions. Merely assuming that the densities are twice continuously differentiable allows for estimation at $O_{p}(n^{-2/(d+4)})$ rate. This rate becomes extremely slow when $d$ is even moderately large. Nagler and Czado, (2016) showed that the simplifying assumption alleviates this. In particular, they construct nonparametric estimators of the simplified model that converge at $O_{p}(n^{-1/3})$ rate — irrespective of the dimension $d$. The simplifying assumption effectively makes estimation as easy as in a bivariate model. This comes at the cost of a potentially large approximation error, but it may be a good trade-off in practice. Nagler and Czado, (2016) show in simulations that, even on large samples, an estimated simplified model can be much closer to the true density than a fully nonparametric estimate, even if the simplifying assumption is violated severely. The gains in estimability simply outweigh the loss in approximation ability of the simplified model.

### v2

@@ Abstract
Vine copula models have become highly popular practical tools for modeling multivariate dependencies. To maintain tractability, a commonly employed simplifying assumption is that conditional copulas remain unchanged by the conditioning variables. This assumption has sparked a somewhat polarizing debate within the copula community. The fact that much of this dispute occurs outside the public record has placed the field in an unfortunate position, impeding scientific progress. In this article, I will review what we know about the flexibility and limitations of simplified vine copula models, explore the broader implications, and offer my own, hopefully reconciling, perspective on the issue.
@@ 2.2 The simplifying assumption in the bivariate case
Now consider the following conditional version of Sklar’s theorem (Patton,, 2000). For any index set $D\subset\{1,\ldots,d\}$, indices $i,j\notin D$, and every value $\bm{x}_{D}\in\mathds{R}^{|D|}$, the conditional distribution of $(X_{i},X_{j})$ given $\bm{X}_{D}=\bm{x}_{D}$ can be expressed as
@@
where $C_{i,j|D}(\cdot\mid\bm{x}_{D})$ is a bivariate copula for every value of $\bm{x}_{D}$. [^1] In the vine copula literature, the conditional copula is more commonly denoted by $C_{i,j;D}$ to distinguish it from the conditional distribution of $(U_{i},U_{j})$ given $\bm{U}_{D}$, which is not a copula. The latter distribution or related conditionals do not appear in this article, so there should be no confusion. More precisely, it is the conditional joint distribution function of the random vector $(U_{i|D},U_{j|D})$ given $\bm{X}_{D}=\bm{x}_{D}$, where
@@
The Gaussian distribution provides an accessible analogy. Let $\bm{X}\sim\mathcal{N}(0,\Sigma)$ and define $\varepsilon_{i|D}$ and $\varepsilon_{j|D}$ as the erorrs from regressing $X_{i}$ and $X_{j}$ on $\bm{X}_{D}$, scaled to unit variance. This is, in essence, what the operation $F_{i|D}(X_{i}\mid\bm{X}_{D})$ does. The conditional copula $C_{i,j|D}(\cdot\mid\bm{x}_{D})$ characterizes the dependence of the regression erorrs $\varepsilon_{i|D}$ and $\varepsilon_{j|D}$. In the Gaussian world, dependence is synonymous with correlation, so we can think of $C_{i,j|D}(\cdot\mid\bm{x}_{D})$ as the conditional correlation coefficient $\rho_{i,j|D}(\bm{x}_{D})={\mathrm{Corr}}(\varepsilon_{i|D},\varepsilon_{j|D}\mid\bm{X}_{D}=\bm{x}_{D})$. The simplifying assumption asserts that this correlation coefficient does not change with the value $\bm{x}_{D}$. Indeed, the Gaussian distribution satisfies $\rho_{i,j|D}\equiv\rho_{i,j;D}$, where $\rho_{i,j;D}$ is the partial correlation coefficient of $(X_{i},X_{j})$ given $\bm{X}_{D}$. This equality does not typically hold outside of Gaussian models, however. For more details, we refer to Gijbels et al., (2015) and Spanhel and Kurz, (2016) who expand on this idea in terms of conditional and partial copulas.
@@ 2.3 The simplifying assumption in vine copulas
3 What we know about the limitations of simplified models
@@ 3.5 Simplified vines as statistical models
Simplified vine copulas are statistical models for the dependence structure of a random vector. From a philosophical point of view, they are no different from other statistical models serving similar purposes. Models are approximations of reality, and as such, mostly wrong. To paraphrase a famous quote by George Box: The question is not whether they are wrong, but whether they are useful.
@@
The second point emphasizes that we cannot disentangle the approximation qualities of a model from our ability to estimate it. Generally speaking, more flexible models are harder to estimate—the ubiquitous bias-variance trade-off. In extreme cases, the model can be so large that consistent estimation is impossible. This is the case, for example, if $\Theta$ is the collection of all probability densities on $\mathds{R}^{d}$. Parametric families of densities on the other hand can be estimated at $O_{p}(n^{-1/2})$ rate under mild conditions. Merely assuming that the densities are twice continuously differentiable allows for estimation at $O_{p}(n^{-2/(d+4)})$ rate. This rate becomes extremely slow when $d$ is even moderately large. Nagler and Czado, (2016) showed that the simplifying assumption alleviates this. In particular, they construct nonparametric estimators of the partial vine copula that converge at $O_{p}(n^{-1/3})$ rate — irrespective of the dimension $d$. The simplifying assumption effectively makes estimation as easy as in a bivariate model. This comes at the cost of a potentially large approximation error, but it may be a good trade-off in practice. Nagler and Czado, (2016) show in simulations that, even on large samples, an estimated simplified model can be much closer to the true density than a fully nonparametric estimate, even if the simplifying assumption is violated severely. The gains in estimability simply outweigh the loss in approximation ability of the simplified model.